OpenStudy (anonymous):

1. Write the equation of the line which passes through (5, –2) and is parallel to x = 4. (1 point) 2. Write the equation of the line which passes through (2, 1) and is perpendicular to x = –2. (1 point) 3. Write the equation of the line which passes through (–4, 2) and is parallel to y = –x + 6 in slope-intercept form. (2 point) 4. Write the equation of the line which passes through (2, –3) and is perpendicular to y = 4x + 7 in standard form. (2 point) 5. Using complete sentences, describe one example of a place in your everyday life of parallel lines and one example of perpendicu

OpenStudy (anonymous):

5. Using complete sentences, describe one example of a place in your everyday life of parallel lines and one example of perpendicular lines. You may also provide a digital image or self-created picture. (4 points; 2 points for each example)

OpenStudy (anonymous):

ask your teacher to use complete sentences to explain how he or she would walk up a flight of steps. what is this constant "complete sentence" motif?

OpenStudy (anonymous):

1. x=5 2. y= 1/2x+5 3. y= -x-2 4. y= -1/4-11 5. Perpendicular: Graph paper Parallel: Binder paper

OpenStudy (anonymous):

pancake55556 can you show your work please

OpenStudy (anonymous):

pancake55556 can you show your work please

OpenStudy (anonymous):

ok. 1. if the equation of the line is x= 4, then it is a vertical line through all points with the x value of 4. to make a line parallel to this line, you would make a vertical line through all points with the x value 5, so it would be x=5. 2. when a line is perpendicular to another, the slope will be the opposite reciprocal (meaning it would be negative instead of positive, or vice versa and that the number's numerator and denominator would be switch, so -2 would become 1/2) and to find b, or the y-intercept, you would substitute the x and y values into the equation. so 1=1/2(2)+b, 1=1+b subtract 1 from both sides, making 0=b, making the equation y=1/2x (sorry, i realize i was wrong :P) 3. when a line is parallel to another line, the slope would be the same, therefore the slope of the new line is -1. you would again substitute in the x and y values, so 2=-1(-4)+b 2=4+b, subtract 4 from both sides, -2=b so the equation is y=-x-2

OpenStudy (anonymous):

number 4 is pretty much the same as number 2, but with different numbers and number 5 is just examples of those lines in daily life

OpenStudy (anonymous):

can i please have # 4

OpenStudy (anonymous):

4. the slope would be the opposite reciprocal of 4, so it would be -1/4. -3=-1/4(2)+b, so -3=-.5+b, add .5 to each side so it would be -2.5=b. the equation would be y=-1/4x-2.5. i'm sorry i was wrong again!!

OpenStudy (anonymous):

THANK YOU SOOOO MUCH FOR YOUR HELP You are the BEST

OpenStudy (anonymous):

no problem :D